# Towards the Colmez Conjecture

### v1

@@ 3.2. Yuan–Zhang Decomposition
[…]
@@
$=$ $\frac{-1}{2[K:\mathbb{Q}]}\left(\sum_{\sigma\in H_{K}}\log\lvert\langle\omega^{\sigma}_{\tau},\eta^{\sigma}_{\overline{\tau}}\rangle_{\operatorname{dR}}\rvert-\sum_{p<\infty}\sum_{\sigma\in H_{K}}(v_{p}(\omega^{\sigma}_{\tau})+v_{p}(\eta^{\sigma}_{\overline{\tau}}))\log p\right),$
@@
Note that if $\tau\not\in\Phi$, then $\mathcal{N}(\mathcal{A},\tau)=0$ and so $h(A,\tau)$ is 00 as well.
@@ 3.3. Colmez Decomposition
We review the decomposition of the Faltings heights of CM-abelian varieties $h(\Phi)$ given in [Col93]. Let $p$ be a prime number and let $\sigma\in H_{K}$ and $\tau\in H_{E}$. The projection $\mathcal{O}_{K}\otimes\mathcal{O}_{E}\to\mathcal{O}_{K}$ given by $\sigma$ gives $H^{1}(\mathcal{A})_{\tau}$ the structure of an $\mathcal{O}_{K}$-module. We then define $v_{p,C}(\omega)$ if $\omega\in H^{1}(A)_{\tau}\otimes_{K,\sigma}\overline{\mathbb{Q}_{p}}$ (resp. $\omega\in H^{0}(A,\Omega_{A}^{g})\otimes_{K,\sigma}\overline{\mathbb{Q}_{p}}$) by $v_{p,C}(\omega)=0$ if $\omega$ is a generator of the $\mathcal{O}_{p}$-module $H^{1}(\mathcal{A})_{\tau}\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$ (resp. $H^{0}(\mathcal{A},\Omega_{\mathcal{A}}^{g})\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$) and $v_{p,C}(x\omega)=v_{p}(x)+v_{p,C}(\omega)$ if $x\in\overline{\mathbb{Q}_{p}}$. We use the superscript $v_{p,C}$ to contrast it with the valuation $v_{p,YZ}$ defined in the previous subsection. On $H^{0}(\mathcal{A},\Omega^{g}_{\mathcal{A}})$, this is identical to the valuation given for the Faltings height.
@@ 4. Comparison
These two heights decompositions differ by $\log\lvert d_{E}\rvert$ where $d_{E}$ is the absolute discriminant of $E$. Namely for any CM field $E$, CM-type $\Phi\subset H_{E}$, and $\tau\in\Phi$, we have
$h(\Phi,\tau)=ht(\Phi,\tau)+\frac{1}{2}\log 2\pi+\frac{1}{4g}\log\lvert d_{E}\rvert-\log\mu_{Art}(a_{\Phi,\tau}^{0}).$
@@
[…]
@@
$2h(\Phi,\tau)=\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\left(\log\lvert\langle\omega_{\tau}^{\sigma},\eta_{\overline{\tau}}^{\sigma}\rangle\rvert-\sum_{p<\infty}\left(v_{p,YZ}(\omega_{\tau}^{\sigma})+v_{p,YZ}(\eta_{\overline{\tau}}^{\sigma})\right)\log p\right).$
@@
(1) $\begin{split}2h(\Phi,\tau)=\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\bigg{(}&\log\left\lvert\langle\omega_{\tau}^{\sigma},u_{\sigma}\rangle\right\rvert+\log\left\lvert\langle\eta_{\overline{\tau}}^{\sigma},v_{\sigma}\rangle\right\rvert-\log 2\pi\\ &-\sum_{p<\infty}\log p(v_{p,YZ}(\omega_{\tau}^{\sigma})+v_{p,YZ}(\eta_{\overline{\tau}}^{\sigma}))\bigg{)}\\ \end{split}$
The height given by Colmez is Galois invariant and thus $ht(\Phi,\tau)=ht(\overline{\Phi},\overline{\tau})$. It is also invariant under choice of abelian variety so we will choose $A$ for calculating $ht(\Phi,\tau)$ and $A^{t}$ for calculating $ht(\overline{\Phi},\overline{\tau})$. By [Col93, Lem. II.2.13], for each $\sigma\in H_{K}$ and $\tau\in H_{E}$, there exists a $\beta_{\sigma,\tau}\in K^{\times}$ (and $\gamma_{\sigma,\overline{\tau}}\in K^{\times}$) such that
@@
Putting this in gives
$ht(\Phi,\tau)+ht(\overline{\Phi},\overline{\tau})$
$=\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\bigg{(}$ $\log\lvert\langle\omega_{\tau}^{\sigma},\omega_{\overline{\tau}}^{\sigma},u_{\sigma}\rangle\rvert+\log\lvert\langle\eta_{\overline{\tau}}^{\sigma},\eta_{\tau}^{\sigma},v_{\sigma}\rangle\rvert$
$=\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\bigg{(}$ $\log\left\lvert\langle\omega_{\tau}^{\sigma},u_{\sigma}\rangle\right\rvert+\log\left\lvert\langle\eta_{\overline{\tau}}^{\sigma},v_{\sigma}\rangle\right\rvert-\frac{1}{2}\log\lvert\sigma(\beta_{\sigma,\tau})\rvert-\frac{1}{2}\log\lvert\sigma(\gamma_{\sigma,\overline{\tau}})\rvert$
@@
The same mutatis mutandis holds true for $\sigma(\gamma_{\sigma,\overline{\tau}})$. Using that $\mathfrak{a}_{\sigma}\overline{\mathfrak{b}_{\sigma}}=\overline{\mathfrak{d}_{E}^{-1}}$, we combine everything to get
(2) $\begin{split}ht(\Phi,\tau)+ht(\overline{\Phi},\overline{\tau})&\\ =\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\bigg{(}&\log\left\lvert\langle\omega_{\tau}^{\sigma},u_{\sigma}\rangle\right\rvert+\log\left\lvert\langle\eta_{\overline{\tau}}^{\sigma},v_{\sigma}\rangle\right\rvert\\ &-\sum_{p<\infty}\log p\left(v_{p,C}(\omega_{\tau}^{\sigma})+v_{p,C}(\eta_{\overline{\tau}}^{\sigma})-\frac{v_{p}(\sigma\tau\mathfrak{d}_{E}^{-1})+v_{p}(\sigma\overline{\tau}\mathfrak{d}_{E}^{-1})}{2}\right)\bigg{)}\\ \end{split}$
@@
$(\displaystyle\frac{1}{2[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\sum_{p<\infty}\log p\bigg{(}$$(v_{p,YZ}(\omega_{\tau}^{\sigma})-v_{p,C}(\omega_{\tau}^{\sigma})+v_{p,YZ}(\eta_{\overline{\tau}}^{\sigma})-v_{p,C}(\eta_{\overline{\tau}}^{\sigma})$
$+\frac{v_{p}(\sigma\tau\mathfrak{d}_{E}^{-1})+v_{p}(\sigma\overline{\tau}\mathfrak{d}_{E}^{-1})}{2}\bigg{)}$
We now compute the local terms. Fix $p<\infty$ and $\sigma\in H_{K}$, then we view $\sigma\colon K\to\overline{\mathbb{Q}_{p}}$ which gives a places $v$ of $K$. By [Col93, Lem. II.1.2], the de Rham cohomology $H^{1}_{\operatorname{dR}}(\mathcal{A}/\mathcal{O}_{K_{v}})$ is a free $\mathcal{O}_{K_{v}}\otimes\mathcal{O}_{E}$-module of rank $1$. After localizing, we have
@@
The valuation $v_{p,YZ}$ is given on $\mathcal{O}_{K_{v},\tau}$ under the the projection map $R_{\Phi_{\mathfrak{p}}}\to\mathcal{O}_{K_{v},\tau}$, which is surjective. On the other hand, the valuation $v_{p,C}$ is given by the intersection
@@
Let $\pi\in\mathcal{O}_{K_{v},\tau}$ be such that $v_{p,C}(\pi)=0$. Write $\mathcal{O}_{E_{\mathfrak{p}}}=\mathbb{Z}_{p}[\alpha]$ so that $R_{\Phi_{\mathfrak{p}}}$ is generated as an $\mathcal{O}_{K_{v}}$-module by $\{(\rho(\alpha^{i}))_{\rho\in\Phi_{\mathfrak{p}}}\}_{0\leq i<\lvert\Phi_{\mathfrak{p}}\rvert}$. Then $\pi$ is such that the $\tau$-th row of the inverse of the Vandermonde matrix associated to $\alpha$ and $\Phi_{\mathfrak{p}}$ has minimum valuation $-v_{p}(\pi)$. The $\tau$-th row of the inverse matrix consists of terms involving combinations of symmetric polynomials of the $\rho(\alpha)$ all divided by $\prod_{\rho\in\Phi_{\mathfrak{p}}\backslash\tau}(\rho(\alpha)-\tau(\alpha))$. Thus, we have that
@@
$h(\Phi,\tau)-ht(\Phi,\tau)-\frac{1}{2}\log 2\pi$
$=\frac{1}{2[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\sum_{p<\infty}\left(\log p\left(\frac{v_{p}(\sigma\tau\mathfrak{d}_{E})+v_{p}(\sigma\overline{\tau}\mathfrak{d}_{E})}{2}\right)-\mu_{Art,p}(a_{\Phi,\tau})-\mu_{Art,p}(a_{\overline{\Phi},\overline{\tau}})\right)$
$=\frac{1}{2[E:\mathbb{Q}]}\log\lvert d_{E}\rvert-\mu_{Art,p}(a_{\Phi,\tau}^{0}),$
@@
By Theorems 4.1 and 2.2, it suffices to show that $a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0}$ depends only on $E$. But, this is because a simple computation shows
$a_{\Phi_{1},\tau_{1}}^{0}(g)+a_{\Phi_{2},\tau_{2}}^{0}(g)=\begin{cases}2&\text{if $g|_{E}=\mathrm{Id}$,}\\ 0&\text{if $g|_{E}=c$,}\\ 1&\text{otherwise.}\end{cases}$
Moreover, this leads to another proof of [YZ18, Cor. 2.6] relating the height of nearby CM-types and Faltings heights of CM-abelian varieties.
@@
The extra $\log 2\pi$ term not seen in [YZ18, Cor. 2.6] is because we use a different normalizing factor for the archimedean contribution in the Faltings height.
@@
We write the left hand side in terms of class functions using linearity. A straightforward calculation shows that
$\frac{1}{g2^{g}}\sum_{\Phi}A_{\Phi}(g)=\begin{cases}1&\text{if $g|_{E}=\mathrm{Id}$,}\\ 0&\text{if $g|_{E}=c$,}\\ \frac{1}{2}&\text{otherwise.}\end{cases}=\frac{a_{\Phi_{1},\tau_{1}}^{0}(g)+a_{\Phi_{2},\tau_{2}}^{0}(g)}{2}$
@@
We may rewrite $a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0}$ in terms of Artin characters as $1+\frac{1}{g}\operatorname{Ind}_{E/F}^{E/\mathbb{Q}}\chi_{E/F}$, where $\chi_{E/F}\colon\operatorname{Gal}(E/F)\to\pm 1$ is given by $\chi_{E/F}(c)=-1$. Thus its log conductor is
@@
$\frac{1}{g2^{g}}\sum_{\Phi}h(\Phi)-\frac{h(\Phi_{1},\tau_{1})+h(\Phi_{2},\tau_{2})}{2}=$ $-\frac{1}{2}\log 2\pi-\frac{1}{4g}\log\lvert d_{E}\rvert+\frac{\mu_{Art}(a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0})}{4}$

### v2

@@ 3.2. Yuan–Zhang Decomposition
For each archimedean place of $K$, we use the aforementioned Hermitian norm $\lVert\cdot\rVert$ on the generic fiber of $\Omega(\mathcal{A})_{\tau}\otimes\Omega(\mathcal{A}^{t})_{\overline{\tau}}$, and thus we get a metrized line bundle
@@
$=$ $\frac{-1}{2[K:\mathbb{Q}]}\left(\sum_{\sigma\in H_{K}}\log\lvert\langle\omega^{\sigma}_{\tau},\eta^{\sigma}_{\overline{\tau}}\rangle_{\operatorname{dR}}\rvert-\sum_{\begin{subarray}{c}p<\infty\\ \sigma\in H_{K}\end{subarray}}(v_{p}(\omega^{\sigma}_{\tau})+v_{p}(\eta^{\sigma}_{\overline{\tau}}))\log p\right)$
@@
Note that if $\tau\not\in\Phi$, then $\mathcal{N}(\mathcal{A},\tau)=0$ and so $h(A,\tau)$ is 0 as well.
@@ 3.3. Colmez Decomposition
We review the decomposition of the Faltings heights of CM-abelian varieties $h(\Phi)$ given in [COL93]. Let $p$ be a prime number and let $\sigma\in H_{K}$ and $\tau\in H_{E}$. The projection $\mathcal{O}_{K}\otimes\mathcal{O}_{E}\to\mathcal{O}_{K}$ given by $\sigma$ gives $H^{1}(\mathcal{A})_{\tau}$ the structure of an $\mathcal{O}_{K}$-module. We then define $v_{p,C}(\omega)$ if $\omega\in H^{1}(A)_{\tau}\otimes_{K,\sigma}\overline{\mathbb{Q}_{p}}$ (resp. $\omega\in H^{0}(A,\Omega_{A}^{g})\otimes_{K,\sigma}\overline{\mathbb{Q}_{p}}$) by $v_{p,C}(\omega)=0$ if $\omega$ is a generator of the $\mathcal{O}_{p}$-module $H^{1}(\mathcal{A})_{\tau}\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$ (resp. $H^{0}(\mathcal{A},\Omega_{\mathcal{A}}^{g})\otimes_{\mathcal{O}_{K},\sigma}\mathcal{O}_{p}$) and $v_{p,C}(\alpha\omega)=v_{p}(\alpha)+v_{p,C}(\omega)$ if $\alpha\in\overline{\mathbb{Q}_{p}}$. We use the superscript $v_{p,C}$ to contrast it with the valuation $v_{p,YZ}$ defined in the previous subsection. On $H^{0}(\mathcal{A},\Omega^{g}_{\mathcal{A}})$, this is identical to the valuation given for the Faltings height.
@@ 4. Comparison
These two heights decompositions differ by $\log\lvert d_{E}\rvert$ where $d_{E}$ is the absolute discriminant of $E$. Namely, for any CM field $E$, CM-type $\Phi\subset H_{E}$, and $\tau\in\Phi$, we have
$h(\Phi,\tau)=ht(\Phi,\tau)+\frac{1}{2}\log 2\pi+\frac{1}{4g}\log\lvert d_{E}\rvert-\mu_{Art}(a_{\Phi,\tau}^{0}).$
@@
is a perfect pairing between rank $2g$ $\mathbb{Z}$-modules. Moreover, both admit a compatible $\mathcal{O}_{E}$-action so that we may identify each with fractional ideals of $E$ and so that the pairing becomes a twist of the trace pairing $\langle u,v\rangle=2\pi i\operatorname{Tr}_{E/\mathbb{Q}}(u\overline{v})$.
For each $\sigma\in H_{K}$, choose nonzero elements $u_{\sigma}\in H_{1}(A^{\sigma}(\mathbb{C}),\mathbb{Z})$ and $v_{\sigma}\in H_{1}(A^{t,\sigma}(\mathbb{C}),\mathbb{Z})$ such that $u_{\overline{\sigma}}=c(u_{\sigma})$ and both are equal to $1$ under our identification of $H_{1}(A^{\sigma},\mathbb{Z})$ and $H_{1}(A^{t,\sigma},\mathbb{Z})$ with fractional ideals of $E$. We can find elements $\alpha_{i,\sigma},\beta_{i,\sigma}\in E$ that form a dual basis of $H_{1}(A^{\sigma},\mathbb{Z})$ and $H_{1}(A^{t,\sigma},\mathbb{Z})$ so that $\langle\alpha_{i,\sigma}u_{\sigma},\beta_{j,\sigma}v_{\sigma}\rangle=2\pi i\delta_{ij}$. Let $\mathfrak{a}_{\sigma}$ denote the fractional ideal generated by the $\alpha_{i,\sigma}$ and $\mathfrak{b}_{\sigma}$ the fractional ideal generated by the $\beta_{i,\sigma}$. Since the pairing is the trace pairing, we have that $\mathfrak{b}_{\sigma}=\overline{\mathfrak{a}_{\sigma}^{-1}\mathfrak{d}_{E}^{-1}}$, where $\mathfrak{d}_{E}$ is the different ideal.
@@
$2h(\Phi,\tau)=\frac{-1}{[K:\mathbb{Q}]}\sum_{\sigma\in H_{K}}\Bigg($ $\log\lvert\langle\omega_{\tau}^{\sigma},\eta_{\overline{\tau}}^{\sigma}\rangle\rvert$
$-\sum_{p<\infty}\left(v_{p,YZ}(\omega_{\tau}^{\sigma})+v_{p,YZ}(\eta_{\overline{\tau}}^{\sigma})\right)\log p\Bigg).$
@@
[…]
The height given by Colmez is Galois invariant and thus $ht(\Phi,\tau)=ht(\overline{\Phi},\overline{\tau})$. It is also invariant under the choice of abelian variety so we will choose $A$ for calculating $ht(\Phi,\tau)$ and $A^{t}$ for calculating $ht(\overline{\Phi},\overline{\tau})$. By [COL93, Lem. II.2.13], for each $\sigma\in H_{K}$ and $\tau\in H_{E}$, there exists a $\beta_{\sigma,\tau}\in K^{\times}$ (and $\gamma_{\sigma,\overline{\tau}}\in K^{\times}$) such that
@@
Putting this in gives that $ht(\Phi,\tau)+ht(\overline{\Phi},\overline{\tau})$ is equal to
[…]
@@
The same, mutatis mutandis, holds true for $\sigma(\gamma_{\sigma,\overline{\tau}})$. Using that $\mathfrak{a}_{\sigma}\overline{\mathfrak{b}_{\sigma}}=\overline{\mathfrak{d}_{E}^{-1}}$, we combine everything to get
[…]
@@
$\frac{1}{2[K:\mathbb{Q}]}\sum_{\begin{subarray}{c}\sigma\in H_{K}\\ p<\infty\end{subarray}}\log p\bigg($ $v_{p,YZ}(\omega_{\tau}^{\sigma})-v_{p,C}(\omega_{\tau}^{\sigma})+v_{p,YZ}(\eta_{\overline{\tau}}^{\sigma})$
$-v_{p,C}(\eta_{\overline{\tau}}^{\sigma})+\frac{v_{p}(\sigma\tau\mathfrak{d}_{E}^{-1})+v_{p}(\sigma\overline{\tau}\mathfrak{d}_{E}^{-1})}{2}\bigg)$
We now compute the local terms. Fix $p<\infty$ and $\sigma\in H_{K}$, then we view $\sigma\colon K\to\overline{\mathbb{Q}_{p}}$ which gives a place $v$ of $K$. By [COL93, Lem. II.1.2], the de Rham cohomology $H^{1}_{\operatorname{dR}}(\mathcal{A}/\mathcal{O}_{K_{v}})$ is a free $\mathcal{O}_{K_{v}}\otimes\mathcal{O}_{E}$-module of rank $1$. After localizing, we have
@@
The valuation $v_{p,YZ}$ is given on $\mathcal{O}_{K_{v},\tau}$ under the projection map $R_{\Phi_{\mathfrak{p}}}\to\mathcal{O}_{K_{v},\tau}$, which is surjective. On the other hand, the valuation $v_{p,C}$ is given by the intersection
@@
Let $\pi\in\mathcal{O}_{K_{v},\tau}$ be such that $v_{p,C}(\pi)=0$. Write $\mathcal{O}_{E_{\mathfrak{p}}}=\mathbb{Z}_{p}[\alpha]$ so that $R_{\Phi_{\mathfrak{p}}}$ is generated as an $\mathcal{O}_{K_{v}}$-module by $\{(\rho(\alpha^{i}))_{\rho\in\Phi_{\mathfrak{p}}}\}_{0\leq i<\lvert\Phi_{\mathfrak{p}}\rvert}$. Then $\pi$ is such that the $\tau$-th row of the inverse of the Vandermonde matrix associated to $\alpha$ and $\Phi_{\mathfrak{p}}$ has minimum valuation $v_{p}(\pi)$. The $\tau$-th row of the inverse matrix consists of terms involving combinations of symmetric polynomials of the $\rho(\alpha)$ all divided by the product
$\prod_{\rho\in\Phi_{\mathfrak{p}}\backslash\tau}(\rho(\alpha)-\tau(\alpha)).$
Thus, we have that
@@
$h(\Phi,\tau)-ht(\Phi,\tau)=$ $\frac{1}{2[K:\mathbb{Q}]}\sum_{\begin{subarray}{c}\sigma\in H_{K}\\ p<\infty\end{subarray}}\bigg(\log p\left(\frac{v_{p}(\sigma\tau\mathfrak{d}_{E})+v_{p}(\sigma\overline{\tau}\mathfrak{d}_{E})}{2}\right)$
$-\mu_{Art,p}(a_{\Phi,\tau})-\mu_{Art,p}(a_{\overline{\Phi},\overline{\tau}})\bigg)+\frac{1}{2}\log 2\pi$
$=$ $\frac{1}{2[E:\mathbb{Q}]}\log\lvert d_{E}\rvert-\mu_{Art,p}(a_{\Phi,\tau}^{0})+\frac{1}{2}\log 2\pi,$
@@
By Theorems 4.1 and 2.2, it suffices to show that $a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0}$ depends only on $E$. But, this is because a simple computation shows that for $x\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$,
$a_{\Phi_{1},\tau_{1}}^{0}(x)+a_{\Phi_{2},\tau_{2}}^{0}(x)=\begin{cases}2&\text{if $x|_{E}=\mathrm{Id}$,}\\ 0&\text{if $x|_{E}=c$,}\\ 1&\text{otherwise.}\end{cases}$
Moreover, this leads to another proof of [YZ18, Cor. 2.6] relating the height of nearby CM-types and Faltings heights of CM abelian varieties.
@@
The extra $\log 2\pi$ term not seen in [YZ18, Cor. 2.6] is because we use a different normalizing factor for the archimedean contribution in the Faltings height. This is expanded on in the remark after Definition 3.1.
@@
We write the left hand side in terms of class functions using linearity. A straightforward calculation shows that for $x\in\operatorname{Gal}(\overline{\mathbb{Q}}/\mathbb{Q})$,
$\frac{1}{g2^{g}}\sum_{\Phi}A_{\Phi}(x)=\frac{a_{\Phi_{1},\tau_{1}}^{0}(x)+a_{\Phi_{2},\tau_{2}}^{0}(x)}{2}=\begin{cases}1&\text{if $x|_{E}=\mathrm{Id}$,}\\ 0&\text{if $x|_{E}=c$,}\\ \frac{1}{2}&\text{otherwise.}\end{cases}$
@@
We have the equality $a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0}=1+\frac{1}{g}\operatorname{Ind}_{E/F}^{E/\mathbb{Q}}\chi_{E/F}$, where $\chi_{E/F}$ is the nontrivial character of $\operatorname{Gal}(E/F)$. Thus its log conductor is
@@
$\frac{1}{g2^{g}}\sum_{\Phi}h(\Phi)-\frac{h(\Phi_{1},\tau_{1})+h(\Phi_{2},\tau_{2})}{2}=$ $-\frac{1}{2}\log 2\pi-\frac{1}{4g}\log\lvert d_{E}\rvert$
$+\frac{\mu_{Art}(a_{\Phi_{1},\tau_{1}}^{0}+a_{\Phi_{2},\tau_{2}}^{0})}{4}$